# Random dynamics and invariant measures for a class of non-Newtonian fluids of differential type on 2D and 3D Poincaré domains

### v1

@@ 5. Random attractors for stochastic third-grade fluids equations: Unbounded domains
$\quad+C\bigg{\{}\|\mathrm{A}(\boldsymbol{y})\|_{2}+\|\boldsymbol{z}\|_{2}\bigg{\}}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}}+\frac{C}{k}\bigg{[}\|\boldsymbol{y}\|^{4}_{2}+\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}$
$\qquad+\|\boldsymbol{z}\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{z}\|^{2}_{2}+\|\boldsymbol{f}\|^{2}_{2}\bigg{]}.$ (5.40)
@@
×$\qquad\times\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}(x,\tau)|^{4}\mathrm{d}x\right)^{\frac{1}{4}}\mathrm{d}\tau$
@@
$\to 0\text{ as }k\to\infty,$ (5.41)
where we have used the finiteness of integral obtained in (5.2)-(5.10) and Lemmas 3.5-3.6. Hence, from (5), we conclude that for any $\varepsilon>0$, for any $\omega\in\Omega$ and for any $\xi\in[-1,0]$, there exists a $k_{0}=k_{0}(\varepsilon,\xi,\omega)\in\mathbb{N}$ such that
@@
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Consider the MDS, $\Im=(\Omega,\mathscr{F},\mathbb{P},\theta)$ from Proposition 3.1, and the RDS $\Psi$ on $\mathbb{H}$ over $\Im$ generated by third-grade fluids equations (2.18) with additive noise satisfying Hypothesis 1.1. Then, there exists a unique random $\mathfrak{DK}$-attractor for continuous RDS $\Psi$ in $\mathbb{H}$.
@@
Because of [13, Theorem 2.8], it is only needed to prove that there exists a $\mathfrak{DK}$-absorbing set $\widetilde{\textbf{B}}\in\mathfrak{DK}$ and the RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. Existence of $\mathfrak{DK}$-absorbing set $\widetilde{\textbf{B}}\in\mathfrak{DK}$: Let $\mathrm{D}$ be a random set from the class $\mathfrak{DK}$. Let $\kappa_{\mathrm{D}}(\omega)$ be the radius of $\mathrm{D}(\omega)$, that is, $\kappa_{\mathrm{D}}(\omega):=\sup\{\|x\|_{2}:x\in\mathrm{D}(\omega)\}$ for $\omega\in\Omega$.
@@
The above inequality implies that for $\omega\in\Omega$, $\boldsymbol{v}(0,\omega;s,\boldsymbol{x})\in\widetilde{\textbf{B}}(\omega)$, for all $s\leq-t_{\mathrm{D}}(\omega).$ This proves $\widetilde{\textbf{B}}$ absorbs $\mathrm{D}$. The RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. In order to establish the $\mathfrak{DK}$-asymptotically compactness of $\Psi$, we use uniform-tail estimates obtained in Lemma 5.4. Let us assume that $\mathrm{D}\in\mathfrak{DK}$ and $\widetilde{\textbf{B}}\in\mathfrak{DK}$ be such that $\widetilde{\textbf{B}}$ absorbs $\mathrm{D}$. Let us fix $\omega\in\Omega$ and take a sequence of positive numbers $\{t_{m}\}^{\infty}_{m=1}$ such that $t_{1}\leq t_{2}\leq t_{3}\leq\cdots$ and $t_{m}\to\infty$. We take an $\mathbb{H}$-valued sequence $\{\boldsymbol{x}_{m}\}^{\infty}_{m=1}$ such that $\boldsymbol{x}_{m}\in\mathrm{D}(\theta_{-t_{m}}\omega),$ for all $m\in\mathbb{N}.$ Our aim is to show that the sequence $\Psi(t_{m},\theta_{-t_{m}}\omega,\boldsymbol{x}_{m})$ or $\boldsymbol{y}(0,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))$ of solutions to system (3.19) has a convergent subsequence in $\mathbb{H}$.
@@
where $C$ is a positive constant independent of $s$ and $\boldsymbol{x}.$ Since $t_{m}\to\infty$, there exists $M_{2}=M_{2}(\omega,\mathrm{D})\in\mathbb{N}$ such that $t_{m}\geq t_{\mathrm{D}}(\omega),$ for all $m\geq M_{2}$. Since $\boldsymbol{x}_{m}\in\mathrm{D}(\theta_{-t_{m}}\omega)$, (5.44) implies that for all $m\geq M_{2}$, the sequence
@@
$\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))=\boldsymbol{y}(\xi,\omega;-1,\boldsymbol{y}(-1,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))-\boldsymbol{z}(-1))\to\hat{\boldsymbol{y}}\ \text{ in }\ \mathbb{L}^{2}(\mathscr{O}_{k}).$ (5.46)
as $m\to\infty$. Therefore, we infer from the proof of Lemma 3.9 that there exists a positive constant $C_{Lip}$ such that
@@ 6. Invariant measures
In this section, we first show the existence of invariant measures for third-grade fluids equations in $\mathbb{H}$. It is established in [27] that the existence of a compact invariant random set is a sufficient condition for the existence of invariant measures, that is, if a RDS $\Psi$ has compact invariant random set, then there exists an invariant measure for $\Psi$ ([27, Corollary 4.4]). Since, the random attractor itself is a compact invariant random set, the existence of invariant measures for the third-grade fluids equations (2.18) is a direct consequence of [27, Corollary 4.4] and Theorem 4.3. The uniqueness of invariant measures is still an open problem.
@@
for all $f\in\mathscr{B}_{b}(\mathbb{H})$, where $\mathscr{B}_{b}(\mathbb{H})$ is the space of all bounded and Borel measurable functions on $\mathbb{H}$ and $\Psi$ is the RDS corresponding to system (2.18), which is defined by (4.1). The continuity of $\Psi$ (cf. Lemma 3.9), [12, Proposition 3.8] provides the following result:
@@
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Then, the family $\{\mathrm{T}_{t}\}_{t\geq 0}$ is Feller, that is, $\mathrm{T}_{t}f\in\mathrm{C}_{b}(\mathbb{H})$ if $f\in\mathrm{C}_{b}(\mathbb{H})$, where $\mathrm{C}_{b}(\mathbb{H})$ is the space of all bounded and continuous functions on $\mathbb{H}$. Furthermore, for any $f\in\mathrm{C}_{b}(\mathbb{H})$, $\mathrm{T}_{t}f(\boldsymbol{x})\to f(\boldsymbol{x})$ as $t\downarrow 0$.
Analogously as in the proof of [27, Theorem 5.6], one can prove that $\Psi$ is a Markov RDS, that is, $\mathrm{T}_{t_{1}+t_{2}}=\mathrm{T}_{t_{1}}\mathrm{T}_{t_{2}}$, for all $t_{1},t_{2}\geq 0$. Since, we know by [27, Corollary 4.4] that if a Markov RDS on a Polish space has an invariant compact random set, then there exists a Feller invariant probability measure $\upmu$ for $\Psi$.
@@
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Then, there exists an invariant measure for system (2.18) in $\mathbb{H}$ with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$ on bounded and unbounded domains, respectively.
Acknowledgments: “This work is funded by national funds through the FCT – Fundação para a Ciência e a Tecnologia, I.P., under the scope of the projects UIDB/00297/2020 (https://doi.org/10.54499/UIDB/ 00297/2020) and UIDP/00297/2020 (https://doi.org/10.54499/UIDP/00297/2020) (Center for Mathematics and Applications)”. K. Kinra would like to thank Prof. Manil T. Mohan, Department of Mathematics, Indian Institute of Technology Roorkee, Roorkee, India for introducing him into the research area of attractor theory.

### v2

@@ 5. Random attractors for a subclass of stochastic third-grade fluids equations: Unbounded domains
$\quad+C\bigg\{\|\mathrm{A}(\boldsymbol{y})\|_{2}+\|\boldsymbol{z}\|_{2}\bigg\}\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|_{4}\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}|^{4}\mathrm{d}x\right)^{\frac{1}{4}}$
$\quad+C\|\nabla\widehat{\textbf{P}}_{1}\|_{\mathbb{H}^{-1}}\bigg(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}|^{2}\mathrm{d}x+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\mathrm{A}(\boldsymbol{y})|^{2}\mathrm{d}x\bigg)^{\frac{1}{2}}$
$\quad+C\|\nabla\widehat{\textbf{P}}_{2}\|_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg(\bigg\{\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}|^{2}\mathrm{d}x\bigg\}^{2}+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{4}\mathrm{d}x+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}|^{4}\mathrm{d}x\bigg)^{\frac{1}{4}}$
$\quad+\frac{C}{k}\bigg[\|\boldsymbol{y}\|^{4}_{2}+\|\mathrm{A}(\boldsymbol{y})\|^{2}_{2}+\|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})\|^{4}_{4}+\|\boldsymbol{z}\|^{4}_{\mathbb{W}^{1,4}}+\|\boldsymbol{z}\|^{2}_{2}+\|\boldsymbol{f}\|^{2}_{2}\bigg].$ (5.40)
@@
$\qquad\times\left(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{z}(x,\tau)|^{4}\mathrm{d}x\right)^{\frac{1}{4}}\mathrm{d}\tau$
$\quad+C\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\|\nabla\widehat{\textbf{P}}_{1}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|_{\mathbb{H}^{-1}}\bigg(\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))|^{2}\mathrm{d}x$
$\qquad+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s)))|^{2}\mathrm{d}x\bigg)^{\frac{1}{2}}\mathrm{d}\tau$
$\quad+C\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\|\nabla\widehat{\textbf{P}}_{2}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))\|_{\mathbb{W}^{-1,\frac{4}{3}}}\bigg(\bigg\{\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))|^{2}\mathrm{d}x\bigg\}^{2}$
$\qquad+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))+\boldsymbol{z}(\tau)|^{4}\mathrm{d}x+\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}(\tau)|^{4}\mathrm{d}x\bigg)^{\frac{1}{4}}\mathrm{d}\tau$
@@
$\quad+C\bigg(\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))|^{2}\mathrm{d}x\mathrm{d}\tau$
$\qquad+\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s)))|^{2}\mathrm{d}x\mathrm{d}\tau\bigg)^{\frac{1}{2}}$
$\quad+C\bigg(\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\bigg\{\int_{\mathscr{O}\cap\{|x|\geq k\}}|\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))|^{2}\mathrm{d}x\bigg\}^{2}\mathrm{d}\tau$
$\qquad+\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\mathrm{A}(\boldsymbol{y}(\tau,\omega;s,\boldsymbol{x}-\boldsymbol{z}(s))+\boldsymbol{z}(\tau)|^{4}\mathrm{d}x\mathrm{d}\tau$
$\qquad+\int_{s}^{0}e^{\nu\lambda\left(1+\frac{\varepsilon_{0}}{2}\right)\tau}\int_{\mathscr{O}\cap\{|x|\geq k\}}|\nabla\boldsymbol{z}(\tau)|^{4}\mathrm{d}x\mathrm{d}\tau\bigg)^{\frac{1}{4}}$
$\rightarrow 0\text{ as }k\rightarrow\infty,$ (5.41)
where we have used the finiteness of integral obtained in (5.2)-(5.10), Lemmas 3.5-3.6 and (A.5) (see Remark A.1). Hence, from (5), we conclude that for any $\varepsilon>0$, for any $\omega\in\Omega$ and for any $\xi\in[-1,0]$, there exists a $k_{0}=k_{0}(\varepsilon,\xi,\omega)\in\mathbb{N}$ such that
@@
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Consider the MDS, $\Im=(\Omega,\mathscr{F},\mathbb{P},\theta)$ from Proposition 3.1, and the RDS $\Psi$ on $\mathbb{H}$ over $\Im$ generated by system (2.30) with additive noise satisfying Hypothesis 1.1. Then, there exists a unique random $\mathfrak{DK}$-attractor for continuous RDS $\Psi$ in $\mathbb{H}$.
@@
Because of [10, Theorem 2.8], it is only needed to prove that there exists a $\mathfrak{DK}$-absorbing set $\widetilde{\textbf{B}}\in\mathfrak{DK}$ and the RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact.
Existence of $\mathfrak{DK}$-absorbing set $\widetilde{\textbf{B}}\in\mathfrak{DK}$: Let $\mathrm{D}$ be a random set from the class $\mathfrak{DK}$. Let $\kappa_{\mathrm{D}}(\omega)$ be the radius of $\mathrm{D}(\omega)$, that is, $\kappa_{\mathrm{D}}(\omega):=\sup\{\|x\|_{2}:x\in\mathrm{D}(\omega)\}$ for $\omega\in\Omega$.
@@
The above inequality implies that for $\omega\in\Omega$, $\boldsymbol{v}(0,\omega;s,\boldsymbol{x})\in\widetilde{\textbf{B}}(\omega)$, for all $s\leq-t_{\mathrm{D}}(\omega).$ This proves $\widetilde{\textbf{B}}$ absorbs $\mathrm{D}$.
The RDS $\Psi$ is $\mathfrak{DK}$-asymptotically compact. In order to establish the $\mathfrak{DK}$-asymptotically compactness of $\Psi$, we use uniform-tail estimates obtained in Lemma 5.4. Let us assume that $\mathrm{D}\in\mathfrak{DK}$ and $\widetilde{\textbf{B}}\in\mathfrak{DK}$ be such that $\widetilde{\textbf{B}}$ absorbs $\mathrm{D}$. Let us fix $\omega\in\Omega$ and take a sequence of positive numbers $\{t_{m}\}^{\infty}_{m=1}$ such that $t_{1}\leq t_{2}\leq t_{3}\leq\cdots$ and $t_{m}\rightarrow\infty$. We take an $\mathbb{H}$-valued sequence $\{\boldsymbol{x}_{m}\}^{\infty}_{m=1}$ such that $\boldsymbol{x}_{m}\in\mathrm{D}(\theta_{-t_{m}}\omega),$ for all $m\in\mathbb{N}.$ Our aim is to show that the sequence $\Psi(t_{m},\theta_{-t_{m}}\omega,\boldsymbol{x}_{m})$ or $\boldsymbol{y}(0,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))$ of solutions to system (3.19) has a convergent subsequence in $\mathbb{H}$.
@@
where $C$ is a positive constant independent of $s$ and $\boldsymbol{x}.$ Since $t_{m}\rightarrow\infty$, there exists $M_{2}=M_{2}(\omega,\mathrm{D})\in\mathbb{N}$ such that $t_{m}\geq t_{\mathrm{D}}(\omega),$ for all $m\geq M_{2}$. Since $\boldsymbol{x}_{m}\in\mathrm{D}(\theta_{-t_{m}}\omega)$, (5.44) implies that for all $m\geq M_{2}$, the sequence
@@
$\boldsymbol{y}(\xi,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))=\boldsymbol{y}(\xi,\omega;-1,\boldsymbol{y}(-1,\omega;-t_{m},\boldsymbol{x}_{m}-\boldsymbol{z}(-t_{m}))-\boldsymbol{z}(-1))\rightarrow\hat{\boldsymbol{y}}\ \text{ in }\ \mathbb{L}^{2}(\mathscr{O}_{k}).$ (5.46)
as $m\rightarrow\infty$. Therefore, we infer from the proof of Lemma 3.9 that there exists a positive constant $C_{Lip}$ such that
@@ 6. Invariant measures
In this section, we first show the existence of invariant measures for a subclass of third-grade fluids equations in $\mathbb{H}$. It is established in [28] that the existence of a compact invariant random set is a sufficient condition for the existence of invariant measures, that is, if a RDS $\Psi$ has compact invariant random set, then there exists an invariant measure for $\Psi$ ([28, Corollary 4.4]). Since, the random attractor itself is a compact invariant random set, the existence of invariant measures for system (2.30) is a direct consequence of [28, Corollary 4.4] and Theorem 4.3. The uniqueness of invariant measures is still an open problem.
@@
for all $f\in\mathscr{B}_{b}(\mathbb{H})$, where $\mathscr{B}_{b}(\mathbb{H})$ is the space of all bounded and Borel measurable functions on $\mathbb{H}$ and $\Psi$ is the RDS corresponding to system (2.30), which is defined by (4.1). The continuity of $\Psi$ (cf. Lemma 3.9), [13, Proposition 3.8] provides the following result:
@@
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Then, the family $\{\mathrm{T}_{t}\}_{t\geq 0}$ is Feller, that is, $\mathrm{T}_{t}f\in\mathrm{C}_{b}(\mathbb{H})$ if $f\in\mathrm{C}_{b}(\mathbb{H})$, where $\mathrm{C}_{b}(\mathbb{H})$ is the space of all bounded and continuous functions on $\mathbb{H}$. Furthermore, for any $f\in\mathrm{C}_{b}(\mathbb{H})$, $\mathrm{T}_{t}f(\boldsymbol{x})\rightarrow f(\boldsymbol{x})$ as $t\downarrow 0$.
Analogously as in the proof of [28, Theorem 5.6], one can prove that $\Psi$ is a Markov RDS, that is, $\mathrm{T}_{t_{1}+t_{2}}=\mathrm{T}_{t_{1}}\mathrm{T}_{t_{2}}$, for all $t_{1},t_{2}\geq 0$. Since, we know by [28, Corollary 4.4] that if a Markov RDS on a Polish space has an invariant compact random set, then there exists a Feller invariant probability measure $\upmu$ for $\Psi$.
@@
Let (1.4) and Hypotheses 1.1-1.2 be satisfied. Then, there exists an invariant measure for system (2.30) in $\mathbb{H}$ with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$ and $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$ on bounded and unbounded domains, respectively.
Remark 6.1.
Note that we discuss the existence of an invariant measure based on the existence of random attractor which is a compact set. But the converse is not true always (see [65]). In particular, the authors in [65] discuss an example of a random dynamical system in $\mathbb{R}^{d}$, $d\geq 2$ namely an isotropic Brownian flow with drift, whose single-point motion is an ergodic diffusion process and which does not have a weak attractor. This also implies that there is no pathwise attractor for such process.
Appendix A Retrieval of Pressure
We have shown that the system (3.19) has unique weak solution in the sense of Definition 3.2 (see Theorem 3.8). In order to retrieve the pressure, we follow the approach which has been carried out in the work [66, Section 8].
[…]
$:=\int_{0}^{T}\biggl\{\bigg\langle\frac{\partial\boldsymbol{y}}{\partial t}-\nu\Delta\boldsymbol{y}+[(\boldsymbol{y}+\boldsymbol{z})\cdot\nabla](\boldsymbol{y}+\boldsymbol{z})-\alpha\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2})-\beta\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))$
$\qquad\qquad-\chi\boldsymbol{z}-\boldsymbol{f},\varphi\bigg\rangle\biggr\}\mathrm{d}t,$ (A.1)
for all $\varphi\in\mathrm{C}_{c}^{\infty}(0,T;\mathrm{C}_{c}^{\infty}(\mathscr{O};\mathbb{R}^{d}))$. Since $\boldsymbol{y}\in\mathrm{C}([0,T];\mathbb{H})\cap\mathrm{L}^{2}(0,T;\mathbb{V})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O})),$ $\boldsymbol{z}\in\mathrm{L}^{2}(0,T;\mathbb{H})\cap\mathrm{L}^{4}(0,T;\mathbb{W}^{1,4}(\mathscr{O}))$ and $\frac{\partial\boldsymbol{y}}{\partial t}\in\mathrm{L}^{2}(0,T;\mathbb{V}^{\prime})+\mathrm{L}^{\frac{4}{3}}(0,T;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O}))$, we have (see e.g. (3.3))
[…]
for all $\varphi\in\mathrm{C}_{c}^{\infty}(0,T;\mathrm{C}_{c}^{\infty}(\mathscr{O};\mathbb{R}^{d}))$. Therefore, it follows that
[…]
Since $\boldsymbol{y}(\cdot)$ is unique solution of system (3.19) in the sense of Definition 3.2, (A) gives
[…]
$=\bigg\langle\frac{\partial\boldsymbol{y}}{\partial t}-\nu\Delta\boldsymbol{y}+[(\boldsymbol{y}+\boldsymbol{z})\cdot\nabla](\boldsymbol{y}+\boldsymbol{z})-\alpha\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2})-\beta\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))-\chi\boldsymbol{z}-\boldsymbol{f},\psi\bigg\rangle,$ (A.4)
for all $\psi\in\mathrm{C}_{c}^{\infty}(\mathscr{O};\mathbb{R}^{d})$, and
[…]
Finally, by an application of [70, Proposition 1.1], there exists a $\hat{\mathbf{P}}\in\mathscr{D}^{\prime}((0,T)\times\mathscr{O})$ such that
[…]
Remark A.1.
Indeed, we have
$\nabla\hat{\mathbf{P}}$ $=(\mathrm{I}-\mathscr{P})\left\{\nu\Delta\boldsymbol{y}-[(\boldsymbol{y}+\boldsymbol{z})\cdot\nabla](\boldsymbol{y}+\boldsymbol{z})+\alpha\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2})+\beta\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))+\boldsymbol{f}\right\}$
$=:\nabla\hat{\mathbf{P}}_{1}+\nabla\hat{\mathbf{P}}_{2},$
where
$\nabla\hat{\mathbf{P}}_{1}$ $=(\mathrm{I}-\mathscr{P})\left\{\nu\Delta\boldsymbol{y}-[(\boldsymbol{y}+\boldsymbol{z})\cdot\nabla](\boldsymbol{y}+\boldsymbol{z})+\alpha\text{div}((\mathrm{A}(\boldsymbol{y}+\boldsymbol{z}))^{2})+\boldsymbol{f}\right\}$
$\nabla\hat{\mathbf{P}}_{2}$ $=\beta(\mathrm{I}-\mathscr{P})\text{div}(|\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})|^{2}\mathrm{A}(\boldsymbol{y}+\boldsymbol{z})).$
Note that, by Lemma 5.2, we have for $s\leq-\mathscr{T}=\max\{t_{\mathrm{D}}^{\ast},t_{\mathrm{D}}^{\ast\ast}\}$ that
$e^{\frac{\nu\lambda}{2}\left(1+\frac{\varepsilon_{0}}{2}\right)\cdot}\;\nabla\widehat{\textbf{P}}_{1}\in\mathrm{L}^{2}(s,0;\mathbb{H}^{-1}(\mathscr{O}))\;\text{ and }\;e^{\frac{3\nu\lambda}{4}\left(1+\frac{\varepsilon_{0}}{2}\right)\cdot}\;\nabla\widehat{\textbf{P}}_{2}\in\mathrm{L}^{\frac{4}{3}}(s,0;\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})).$ (A.5)
Acknowledgments: “This work is funded by national funds through the FCT - Fundação para a Ciência e a Tecnologia, I.P., under the scope of the projects UIDB/00297/2020 (https://doi.org/10.54499/UIDB/00297/2020) and UIDP/00297/2020 (https://doi.org/10.54499/UIDP/00297/2020) (Center for Mathematics and Applications)”. K. Kinra would like to thank Prof. Manil T. Mohan, Department of Mathematics, Indian Institute of Technology Roorkee, Roorkee, India for introducing him into the research area of attractor theory.